OpenStudy (anonymous):

help? Solve for x: 3(x + 1) = -2(x - 1) - 4. 1 -1 -5 -25

OpenStudy (gorv):

open the brackets 3*x+3*1=-2*x-2*(-1)-4

OpenStudy (gorv):

try it now

OpenStudy (triciaal):

3*x+3*1=-2*x-2*(-1)-4 this from above is correct what do you get when you continue?

OpenStudy (anonymous):

ummm i ended up with 0 then -6

OpenStudy (anonymous):

Nevermind , ich werde nur erraten danke trotzdem

OpenStudy (triciaal):

show your work I can find your error you already know the answer is -1

OpenStudy (anonymous):

... um thx buut if i knew i wouldnt have asked thx

OpenStudy (anonymous):